# Dividing Indivisible Items for the Benefit of All: It is Hard to Be Fair Without Social Awareness

### v1

@@ 5 Limited Social Awareness
Some Agents Socially Unaware
@@
For every fixed number of agents n{{n}} and any $c\leq{{n}}$}, there exists a pseudo-polynomial time algorithm deciding whether a SIM allocation, which is simultaneously $\mathcal{F}$-fair for $a_{i}$, $i\in[c]$, and SA-$\mathcal{F}$-fair for any $a_{i}$, $i\in[c+1,\ldots,{{n}}]$, exists for any $\mathcal{F}\in\{\text{tEF1},\text{EF1},\text{sEF1},\text{wEF1},\text{swEF1},\text{EFL}\}$.
Approximate Awareness
@@
We say that an allocation $\mathcal{A}$} is $\alpha$-envy-free up to one item with socially aware agents ($\alpha$-SA-EF1) for $0\leq\alpha\leq 1$, if for each pair of agents $i,j\in{{N}}$, one of the following conditions hold: 1/ there is no EF1-envy from $i$ towards $j$, or 2/ $s_{i}({A}_{j})<\alpha\cdot s_{j}({A}_{j})$.
@@
For any $\alpha\in\mathbb{R}$, $0\leq\alpha<1$, SIM and $\alpha$-SA-EF1 is not guaranteed to exist.
@@
Let $\alpha$ be as stated. We construct an instance with two agents $a_{1}$ and $a_{2}$ and two items $g_{1}$ and $g_{2}$. The valuation functions are identical and assign the same value of $1$ to both items. The social impact functions are $s_{a_{1}}(g_{j})=1$ and $s_{a_{2}}(g_{j})=\alpha$. Due to the social impact function, both items have to be allocated to agent $a_{1}$. Moreover, ${v}_{a_{2}}({A}_{1})=2$ and ${v}_{a_{2}}({A}_{2})=0$}, meaning that agent $a_{2}$ is EF1-envious towards $a_{1}$. The social impacts are $s_{a_{1}}({A}_{1})=2$ and $s_{a_{2}}({A}_{1})=2\alpha$. That is, the second condition from the definition of $\alpha$-SA-EF1 is not met ∎
@@
For any $\alpha\in\mathbb{R}$, $0\leq\alpha<1$, it is NP-complete to decide whether there is a SIM and $\alpha$-SA-EF1 allocation, even if there are only two agents.
Both Proposition 2 and Theorem 11 hold also in the setting where only agent $a_{1}$ is not fully socially aware, which paints our results even stronger.
Weak Social Awareness
@@
Let us have two agents $1$ and $2$, three items $g_{1}$, $g_{2}$, $g_{3}$, and let the valuations and social impacts be as follows.
@@
We say that an allocation $\mathcal{A}$} is envy-free up to one item with weakly socially aware agents (WSA-EF1), if for every pair of agents $i,j\in{{N}}$, one of the following conditions hold: 1/ there is no EF1-envy from $i$ towards $j$, or 2/ ${v}_{i}({A}_{j})\cdot{s}_{i}({A}_{j})\leq{v}_{i}({A}_{i})\cdot s_{j}({A}_{j})$.
The WSA variants for other fairness notions are then defined analogously.
@@
For any $\mathcal{F}\in\{\text{tEF1},\text{sEF1},\text{wEF1},\text{swEF1},\\ \text{EF1},\text{EFL}\}$, it is NP-complete to decide whether SIM and WSA-$\mathcal{F}$-fair allocation exist, even if the social impact function is binary and n=2{{n}}=2.
6 Discussion
We have presented a comprehensive study of computing allocations that maximize social impact while maintaining some fairness criterion. Our results completely resolve the complexity of the problem for goods when the fairness criterion is based on EF1 or EFL.
It is evident from our results that to achieve SIM, social awareness is essential for tractability. An immediate question is whether there exist other natural definitions of social awareness that allow for SIM and fair allocations.

### v2

@@ 5. Limited Social Awareness
5.1. Some Agents Socially Unaware
@@
Proof.
We reduce from the Partition problem. Recall that in this problem, we are given a multi-set of integers $W=(w_{1},\ldots,w_{\ell})$ such that $\sum_{j\in[\ell]}w_{j}=2t$, and the goal is to decide whether $J\subseteq[\ell]$ exists such that $\sum_{j\in J}w_{j}=\sum_{j\in[\ell]\setminus J}w_{j}=t$. This problem is known to be (weakly) NP-hard (GareyJ1979, 31). Without loss of generality, we can assume that no $w_{j}\in W$ is of size $\geq t$, as otherwise we have an obvious Yes (if $w_{j}=t$) or No ($w_{j}>t$) instance.
[…]
 $G_{1},G_{2}$  $G_{3}$  $g_{j}$$j\in[\ell]$
 $a_{1}$  s=0,v=ts=0,{v}=t  s=1,v=ts=1,{v}=t  $s=1,{v}=w_{j}$
 $a_{2}$  s=1,v=0s=1,{v}=0  s=1,v=ts=1,{v}=t  $s=1,{v}=w_{j}$
For the correctness, let $J\subseteq[\ell]$ be a solution for $\mathcal{I}$. We define the bundles as follows. We set ${A}_{1}=\{G_{3}\}\cup\{g_{i}\mid i\in J\}$ and ${A}_{2}=\{G_{1},G_{2}\}\cup\{g_{i}\mid i\in[\ell]\setminus J\}$. Recall that agent $a_{1}$ is not socially aware and ${v}_{a_{1}}({A}_{1})=2t$, while ${v}_{a_{1}}({A}_{2})=3t$. That is, agent $a_{1}$ is envious of $a_{2}$’s bundle; however, this envy can be removed if $G_{1}$ is deleted from $a_{2}$’s bundle. For agent $a_{2}$, on the other hand, we have that ${v}_{a_{2}}({A}_{2})=t$, while ${v}_{a_{2}}({A}_{1})=2t$. Nevertheless, if we remove $G_{3}$ from $a_{1}$’s bundle, we obtain that ${v}_{a_{2}}({A}_{1}\setminus\{G_{3}\})=2t-t=t$. In addition, observe that the allocation indeed maximizes the social impact. That is, the allocation $\mathcal{A}$} is EF1 for both agents.
In the opposite direction, let $\mathcal{A}$} be an allocation which is EF1 for $a_{1}$ and SA-EF1 for $a_{2}$. First we will prove that in any solution $G_{3}\in{A}_{1}$. Then, we will prove that indeed any solution to our problem gives a solution to the original Partition instance. Since we are interested in social impact maximizing allocation, it must hold that $\{G_{1},G_{2}\}\subseteq{A}_{2}$.
For the sake of contradiction assume that  $G_{3}\in{A}_{2}$. Then, agent $a_{1}$ values ${A}_{2}$ at least $3t$. If ${v}_{1}({A}_{2})>3t$, then ${v}_{1}({A}_{2}\setminus\{g\})>2t$ for every $g\in{A}_{2}$. Hence, we cannot eliminate envy for agent 1 in this case. Thus, it must be true that ${v}_{a_{1}}({A}_{2})=3t$. This implies that all remaining items belong to ${A}_{1}$. But in this case we have that ${v}_{a_{2}}({A}_{1})=2t$, thus ${v}_{2}({A}_{1}\setminus\{g\})>t$, for every $g$, and $s_{a_{2}}({A}_{1})=s_{a_{1}}({A}_{1})$. Hence, SA-EF1 criterion is not satisfied for the socially aware agent $a_{2}$. Hence, we conclude that $G_{3}\in{A}_{2}$.
Let now  ${A}_{1}=J\cup\{G_{3}\}$ and ${A}_{2}=\bar{J}\cup\{G_{1},G_{2}\}$, where $(J,\bar{J})$ is a partition of the non-big items. We claim that it must be true that in any acceptable solution it must be true that ${v}_{a_{1}}({A}_{1}\setminus\{G_{3}\})={v}_{a_{1}}({A}_{2}\setminus\{G_{1},G_{2}\})={v}_{a_{2}}({A}_{1}\setminus\{G_{3}\})={v}_{a_{2}}({A}_{2}\setminus\{G_{1},G_{2}\})=t$. In other words, $J$ is a valid solution for the Partition instance. For the sake of contradiction, assume that ${v}_{a_{2}}({A}_{2})={v}_{a_{2}}({A}_{2}\setminus\{G_{1},G_{2}\})=2t-\delta$ for some $\delta>0$. Observe that ${v}_{a_{2}}({A}_{2})={v}_{a_{2}}(\bar{J})$. In addition, observe that
${v}_{a_{2}}({A}_{1})$ $={v}_{a_{2}}(G_{2})+{v}_{a_{2}}(J)$
$=t+2t-{v}_{a_{2}}(\bar{J})$
$>2t+\delta.$
Thus, ${v}_{a_{2}}({A}_{1}\setminus\{g\})>2t>{v}_{a_{2}}({A}_{2})$. In addition, observe that $s_{a_{2}}({A}_{1})=s_{a_{2}}({A}_{1})$. Thus, the conditions of SA-EF1 are not satisifed for agent $a_{2}$. Hence, we conclude that it cannot be true that ${v}_{a_{2}}({A}_{2})<t$, in any solution of our problem. Hence, it must hold that ${v}_{a_{2}}({A}_{2})\geq t$.
Suppose now, again for the sake of contradiction, that ${v}_{a_{2}}({A}_{2})=t+\delta$, for some $\delta>0$. Thus, we get that  ${v}_{a_{2}}({A}_{2}\setminus\{G_{1},G_{2}\})={v}_{a_{2}}(\bar{J})=t+\delta$. Recall, by construction ${v}_{a_{2}}(\bar{J})={v}_{a_{1}}(\bar{J})$ for every set $J$. Then, we get that ${v}_{a_{1}}({A}_{2})={v}_{a_{1}}(\{G_{1},G_{2}\})+{v}_{a_{1}}(\bar{J})=3t+\delta$. Moreover, ${v}_{a_{1}}({A}_{1})={v}_{a_{1}}(G_{3})+{v}_{a_{1}}(J)=2t-\delta$. Hence, ${v}_{a_{1}}({A}_{2}\setminus\{g\})>2t+\delta>{v}_{a_{1}}({A}_{1})$. Hence, the EF1 criterion for agent $a_{1}$ is not satisfied. A contradiction. Hence, we can conclude that ${v}_{a_{2}}({A}_{2})=t$, which in turn means that ${v}_{a_{2}}(\bar{J})={v}_{a_{2}}(J)=t$, i.e., $J$ is a solution for the initial instance of Partition, which completes the proof. ∎
@@
Proof.
If all items can go to both agents, i.e., social impact functions are identical, the question reduces to deciding whether an EF1 allocation exists. It is known that, e.g., by the round-robin mechanism, EF1 allocation always exists and can be found in polynomial time. Otherwise, there is at least one item $g$ in which the socially unaware agent $a_{1}$ is the unique agent that maximizes social impact. In this case, we allocate all items to agent $a_{1}$. Even though the agent $a_{2}$ is very envious of $a_{1}$, it holds that $s_{a_{2}}({A}_{a_{1}})<s_{a_{1}}({A}_{a_{1}})$ because of $g$. Therefore, there is no SA-EF1-envy from $a_{2}$ to $a_{1}$, so such an allocation is indeed a solution. ∎
@@
For every fixed number of agents n{{n}} and any $c\leq{{n}}$}, there exists a pseudo-polynomial time algorithm deciding whether a SIM allocation, which is simultaneously $\mathcal{F}$-fair for $a_{i}$$i\in[c]$, and SA-$\mathcal{F}$-fair for any $a_{i}$$i\in[c+1,\ldots,{{n}}]$, exists for any $\mathcal{F}\in\{\text{tEF1},\text{EF1},\text{sEF1},\text{wEF1},\text{swEF1},\text{EFL}\}$.
Proof.
The algorithm is the same as in Theorem 5, we just add for each (ordered) pair of agents $a$ and $b$ an indication whether $b$ contains an item $g$ such that ${s}_{b}(g)>{s}_{a}(g)$. If this indication is true, then $a$ can no longer envy agent $b$ and we adapt the edges of the graph accordingly. ∎
5.2. Approximate Awareness
@@
We say that an allocation $\mathcal{A}$} is $\alpha$-envy-free up to one item with socially aware agents ($\alpha$-SA-EF1) for $0\leq\alpha\leq 1$, if for each pair of agents $i,j\in{{N}}$, one of the following conditions hold:
  1. (1)
    there is no EF1-envy from $i$ towards $j$, or
  2. (2)
    $s_{i}({A}_{j})<\alpha\cdot s_{j}({A}_{j})$.
@@
For any $\alpha\in\mathbb{R}$$0\leq\alpha<1$, SIM and $\alpha$-SA-EF1 is not guaranteed to exist.
@@
Let $\alpha$ be as stated. We construct an instance with two agents $a_{1}$ and $a_{2}$ and two items $g_{1}$ and $g_{2}$. The valuation functions are identical and assign the same value of $1$ to both items. The social impact functions are
$s_{a_{1}}(g_{j})=1\qquad\text{and}\qquad s_{a_{2}}(g_{j})=\alpha\,.$
Due to the social impact function, both items have to be allocated to agent $a_{1}$. Moreover, ${v}_{a_{2}}({A}_{1})=2$ and ${v}_{a_{2}}({A}_{2})=0$}, meaning that agent $a_{2}$ is EF1-envious towards $a_{1}$. The social impacts are $s_{a_{1}}({A}_{1})=2$ and $s_{a_{2}}({A}_{1})=2\alpha$. That is, the second condition from the definition of $\alpha$-SA-EF1 is not met.∎
@@
For any $\alpha\in\mathbb{R}$$0\leq\alpha<1$, it is NP-complete to decide whether there is a SIM and $\alpha$-SA-EF1 allocation, even if there are only two agents.
Proof.
We reduce from the Partition problem. Recall that in this problem, we are given a multi-set of integers $W=(w_{1},\ldots,w_{\ell})$ such that $\sum_{j\in[\ell]}w_{j}=2t$, and the goal is to decide whether $J\subseteq[\ell]$ exists such that $\sum_{j\in J}w_{j}=\sum_{j\in[\ell]\setminus J}w_{j}=t$. This problem is known to be (weakly) NP-hard (GareyJ1979, 31). Without loss of generality, we can assume that no $w_{j}\in W$ is of size $\geq t$, as otherwise we have an obvious Yes (if $w_{j}=t$) or No ($w_{j}>t$) instance.
Given an instance $\mathcal{I}$ of Partition, we construct the following equivalent instance of our problem as follows. We have two agents $a_{1}$ and $a_{2}$. Next, we have an item $g_{j}$ for every element $w_{j}\in W$. Both agents value this item as $w_{j}$ and are maximizing social impact for these items; that is, $s_{a_{i}}(g_{j})=1$ for every $i\in[2]$. Additionally, we add three “big” items $G_{1}$$G_{2}$, and $G_{3}$. Agent $a_{1}$ values $G_{1},G_{2}$ as 0, while he values $G_{3}$ as $t$. Agent $a_{2}$ values all three big items $G_{1},G_{2},G_{3}$ as $t$. Before we define the social impact, we define
$x=2\cdot\nicefrac{{\ell}}{{\alpha}}\quad\text{and}\quad y=\nicefrac{{\ell}}{{\alpha}}\,.$
Then, the social impact functions are as follows ${s}_{a_{1}}(G_{1})={s}_{a_{1}}(G_{2})=x$${s}_{a_{2}}(G_{1})={s}_{a_{2}}(G_{2})={s}_{a_{1}}(G_{3})=0$, and ${s}_{a_{2}}(G_{3})=y$. The construction is summarized in the following table.
 $G_{1}$$G_{2}$  $G_{3}$  $g_{j}$$j\in[\ell]$
 $a_{1}$  s=x,v=0s=x,{v}=0  s=0,v=ts=0,{v}=t  $s=1,{v}=w_{j}$
 $a_{2}$  s=0,v=0s=0,{v}=0  s=y,v=ts=y,{v}=t  $s=1,{v}=w_{j}$
For the correctness, let $J\subseteq[\ell]$ be a solution for $\mathcal{I}$. We define the bundles as follows. We set ${A}_{1}=\{G_{1},G_{2}\}\cup\{g_{i}\mid i\in J\}$ and ${A}_{2}=\{G_{3}\}\cup\{g_{i}\mid i\in[\ell]\setminus J\}$. Clearly, this allocation is SIM. Now, ${v}_{a_{1}}({A}_{1})=t$, while ${v}_{a_{1}}({A}_{2})=2t$. That is, agent $a_{1}$ is envious of $a_{2}$’s bundle; however, this envy can be removed if $G_{3}$ is deleted from ${A}_{2}$. For agent $a_{2}$, on the other hand, we have that ${v}_{a_{2}}({A}_{2})=2t$, while ${v}_{a_{2}}({A}_{1})=3t$. That is, agent $a_{2}$ is also envious of $a_{1}$. SIf we remove $G_{1}$ from $a_{1}$’s bundle, this envy is eliminated. That is, the allocation $\mathcal{A}$} is $\alpha$-SA-EF1 for both agents.
In the opposite direction, let $\mathcal{A}$} be a MIS allocation. It must hold that $\{G_{1},G_{2}\}\subseteq{A}_{1}$ and $G_{3}\in{A}_{2}$. Now, we construct $J$ so that $J=\{j\in[\ell]\mid g_{j}\in{A}_{1}\}$ and we claim that it is a solution for $\mathcal{I}$. For the sake of contradiction, assume that $\sum_{j\in J}w_{j}\not=\sum_{j\in[\ell]\setminus J}w_{j}$. First, let $\sum_{j\in J}w_{j}>\sum_{j\in[\ell]\setminus J}w_{j}$. Then, ${v}_{a_{2}}({A}_{2})=2t-\delta$ and ${v}_{a_{2}}({A}_{1})=3t+\delta$, where $\delta=|\sum_{j\in[\ell]\setminus J}w_{j}-\sum_{j\in J}w_{j}|$. That is, the agent $a_{2}$ is envious of $a_{1}$, and this envy persists even if we remove $G_{1}$ or $G_{2}$ from ${A}_{1}$. This contradicts that  $\mathcal{A}$} is EF1 for $a_{2}$. Thus, if also $\alpha{s}_{a_{2}}({A}_{2})\leq{s}_{a_{1}}({A}_{2})$, then this cannot be the case. This is equivalent to
$\alpha(y+|[\ell]\setminus J|)$ $\geq|[\ell]\setminus J|$
$\alpha$ $\geq\frac{|[\ell]\setminus J|}{y+|[\ell]\setminus J|}$
and the later holds by the initial setting of $y$, since the nominator is largest if $|[\ell]\setminus J|=\ell$ and the denominator if $|[\ell]\setminus J|=0$. In this case we require that $\alpha\geq\frac{\ell}{y}$. If, on the other hand, we have $\sum_{j\in J}w_{j}<\sum_{j\in[\ell]\setminus J}w_{j}$. Then, ${v}_{a_{1}}({A}_{1})=t-\delta$ and ${v}_{a_{1}}(2t+\delta)$. Again, this envy persists even if we remove $G_{3}$ from ${A}_{2}$. This again contradicts that $\mathcal{A}$} is EF1 for $a_{1}$. Thus, in order for this not to be allowed, we also need that $\alpha{s}_{a_{1}}({A}_{1})\geq{s}+{a_{2}}({A}_{1})$. This is equivalent to
$\alpha(2x+|J|)$ $\geq|J|$
$\alpha$ $\geq\frac{|J|}{2x+|J|}$
and the later holds by the initial setting of $x$, since the nominator is largest if $|J|=\ell$ and the denominator if $|J|=0$. Hence, it must be the case that $\sum_{j\in J}w_{j}=\sum_{j\in[\ell]\setminus J}w_{j}$, which finishes the proof. ∎
Observe that both Proposition 6 and Theorem 7 hold also in the setting where only agent $a_{1}$ is not fully socially aware, which paints our results even stronger.
5.3. Weak Social Awareness
@@
Let us have two agents $1$ and $2$, three items $g_{1}$$g_{2}$$g_{3}$, and let the valuations and social impacts be as follows.
@@
We say that an allocation $\mathcal{A}$} is envy-free up to one item with weakly socially aware agents (WSA-EF1), if for every pair of agents $i,j\in{{N}}$, one of the following conditions hold:
  1. (1)
    there is no EF1-envy from $i$ towards $j$, or
  2. (2)
    ${v}_{i}({A}_{j})\cdot{s}_{i}({A}_{j})\leq{v}_{i}({A}_{i})\cdot s_{j}({A}_{j})$.
The WSA variants for other fairness notions are then defined analogously, with appropriate modifications to condition 1.
@@
For any $\mathcal{F}\in\{\text{tEF1},\text{sEF1},\text{wEF1},\text{swEF1},\text{EF1},\text{EFL}\}$, it is NP-complete to decide whether SIM and WSA-$\mathcal{F}$-fair allocation exist, even if the social impact function is binary and n=2{{n}}=2.
Proof.
This time, we reduce from a variant of the Partition problem called Equitable Partition. In this variant, we are given a multiset $W$ of $2\ell$ integers whose sum is $2t$, and the goal is to decide whether $J\subseteq[2\ell]$ of size $\ell$ exists such that $\sum_{j\in J}w_{j}=\sum_{j\in[2\ell]\setminus J}w_{j}=t$. This problem is known to be NP-hard, even if for each $J\subseteq[2\ell]$ of size less than $\ell$ we have $\sum_{j\in J}w_{j}<t$ (DeligkasEKS2024, 27). Without loss of generality, we can assume that $\ell>4$, as otherwise, we can solve such instance in polynomial by a simple brute-force.
Let $\mathcal{I}$ be an instance of Equitable Partition. We construct an equivalent instance of WSA-EF1-Fair Division with Social Impact as follows (again, we discuss the other notion of fairness at the end of the proof). There are two agents, $a_{1}$ and $a_{2}$. The set of items contains two large items $G_{1}$ and $G_{2}$ together with $2\ell$ small items $g_{j}$$j\in[2\ell]$, which are in one-to-one correspondence with integers of $W$. The valuations and the social impact functions are defined according to the following table.
 $G_{1}$  $G_{2}$  $g_{j}$$j\in[2\ell]$
 $a_{1}$  s=1,v=0s=1,{v}=0  s=0,v=ts=0,{v}=t  $s=1,{v}=w_{j}$
 $a_{2}$  s=0,v=ts=0,{v}=t  s=1,v=0s=1,{v}=0  $s=1,{v}=w_{j}$
For correctness, let $\mathcal{I}$ be a Yes-instance, and let $J$ be a solution. We set ${A}_{a_{1}}=\{G_{1}\}\cup\{g_{j}\mid j\in J\}$ and ${A}_{a_{2}}={M}\setminus{A}_{a_{1}}$. Clearly, the allocation is SIM, since $G_{1}$ is allocated to $a_{1}$ and $G_{2}$ is allocated to $a_{2}$. Now, we claim that $\mathcal{A}$} is WSA-EF1. For $a_{1}$, we have ${v}_{a_{1}}({A}_{a_{2}})=t$ and ${v}_{a_{1}}({A}_{a_{2}})=2t$; however, this envy can be eliminated by removing $G_{2}$ from ${A}_{a_{2}}$. The situation for $a_{2}$ is symmetric: we have ${v}_{a_{2}}({A}_{a_{2}})=t$ and ${v}_{a_{2}}({A}_{a_{1}})=2t$ and envy can be eliminated by removing $G_{1}$ from ${A}_{a_{1}}$.
In the opposite direction, let ${\mathcal{A}}=({A}_{a_{1}},{A}_{a_{2}})$ be a SIM and WSA-EF1 allocation for $\mathcal{J}$. Since $\mathcal{A}$} is SIM, it must be the case that $G_{1}\in{A}_{a_{1}}$ and $G_{2}\in{A}_{a_{2}}$. For the sake of contradiction, assume that $|{A}_{a_{1}}|\not=|{A}_{a_{2}}|$, and, without loss of generality, let $|{A}_{a_{1}}|<|{A}_{a_{2}}|$. Then ${A}_{a_{1}}$ contains at most $\ell-1$ small items. Consequently, ${v}_{a_{1}}({A}_{a_{1}})=t-\delta$ and ${v}_{a_{1}}({A}_{a_{2}})=t+t+\delta=2t+\delta$ and this envy cannot be eliminated even if we remove $G_{2}$ from ${A}_{a_{2}}$. Hence, as $\mathcal{A}$} is WSA-EF1, it must be the case that
${v}_{a_{1}}({A}_{a_{2}})\cdot{s}_{a_{1}}({A}_{a_{2}})\leq{v}_{a_{1}}({A}_{a_{1}})\cdot{s}_{a_{2}}({A}_{a_{2}})\,.$
However, if we substitute the actual values, we obtain
$(2t+\delta)\cdot(|{A}_{a_{2}}|-1)$ $\leq(t-\delta)(|{A}_{a_{2}}|)$
$2t|{A}_{a_{2}}|-2t+\delta|{A}_{a_{2}}|-\delta$ $\leq t|{A}_{a_{2}}|-\delta|{A}_{a_{2}}|$
$t|{A}_{a_{2}}|-2t+2\delta|{A}_{a_{2}}|-\delta$ $\leq 0\,$,
which is clearly never satisfied, as $|{A}_{a_{2}}|>\ell+1$ and $\ell>4$ by our assumptions. Therefore, $\mathcal{A}$} does satisfy neither the second condition of WSA-EF1, which is a contradiction. Hence, it must be the case that $|{A}_{a_{1}}|\geq|{A}_{a_{2}}|$. By symmetric argument, we have that also $|{A}_{a_{2}}|\geq|{A}_{a_{1}}|$, meaning that $|{A}_{a_{1}}|=|{A}_{a_{2}}|=\ell$. We create $J$ such that $J=\{j\in[2\ell]\mid g_{j}\in{A}_{a_{1}}\}$. By the previous arguments, we have $|J|=\ell$. It remains to show that $\sum_{j\in J}w_{j}=t$. Without loss of generality, let $\sum_{j\in J}w_{j}<t$. Then, ${v}_{a_{1}}({A}_{a_{2}})=2t+\delta$ and ${v}_{a_{1}}({A}_{a_{1}})=t-\delta$. However, we have already shown that in this situation, $\mathcal{A}$} is not WSA-EF1, so we have a contradiction. Therefore, it must be the case that $\sum_{j\in J}w_{j}=t$, which finishes the proof. ∎
6. Discussion
We have presented a comprehensive study of computing allocations that maximize social impact while maintaining some fairness criterion. We believe that our results resolve almost completely the problem for goods when the fairness criterion is based on EF1, or EFL. It is evident that in order to achieve SIM, social awareness is essential. Additionally, our results from Section 5 indicate that probably the definition of social awareness from (FlamminiGV2025, 29), i.e., the one we have adopted, is probably the “correct” one, as both relaxations we have identified remove the existential guarantee and furthermore make the problem hard. Hence, an immediate question is whether there exist other natural definitions of social awareness that allow for SIM and fair allocations. In the following, we highlight three other directions that we believe deserve further study.